The Research & Development (R&D) team of a manufacturing firm discovered a new procedure that can be introduced into the production process that can substantially decrease production time. The firm has a limited fund allocated for process modifications that a standard is set for acceptability, that only a production time decrease of at least 8% is acceptable. Six production trials were undertaken showing a mean production time decrease of 8.4% with a standard deviation of 0.32%. Using a level of significance of (a) 0.01 and (b) 0.05, test the hypothesis that the process should be introduced. Use t-distribution
1
Expert's answer
2020-05-25T21:03:23-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\geq8"

"H_1:\mu<8"

This corresponds to a left-tailed test, for which a t-test for one mean, with unknown population standard deviation will be used.

Based on the information provided, the significance level is "\alpha=0.01," and the critical value for a left-tailed test is "t_c=-3.365". The rejection region for this left-tailed test is  "R=\{t:t<-3.365.\}"

The t-statistic is computed as follows:


"t={\bar{X}-\mu\over s/\sqrt{n}}={8.4-8\over 0.32/\sqrt{6}}\approx3.0619"

Since it is observed that "t=3.0619>t_c=-3.365," it is then concluded that the null hypothesis is not rejected. Therefore, there is not enough evidence to claim that the population mean "\mu"  is less than 8, at the 0.01 significance level.

Using the P-value approach: The p-value is "p=0.986," and since "p=0.986>0.01," it is concluded that the null hypothesis is not rejected.


b) Based on the information provided, the significance level is "\alpha=0.05," and the critical value for a left-tailed test is "t_c=-2.015." The rejection region for this left-tailed test is  "R=\{t:t<-2.015\}"

The t-statistic is computed as follows:


"t={\bar{X}-\mu\over s/\sqrt{n}}={8.4-8\over 0.32/\sqrt{6}}\approx3.0619"

Since it is observed that "t=3.0619>-2.015=t_c," it is then concluded that the null hypothesis is not rejected. Therefore, there is not enough evidence to claim that the population mean "\mu"  is less than 8, at the 0.05 significance level.

Using the P-value approach: The p-value is "p=0.986," and since "p=0.986>0.05," it is concluded that the null hypothesis is not rejected.

The process should not be introduced.